3.9(x-2)-1.4x=0.7x+1.38
Answer by b_howie(8)

3.9x-7.8-1.4x=.7x+1.38
2.5x-7.8=.7x+1.38
2.5x-9.18=.7x
-9.18=-1.8x
5.1=x